Great school starts and so does my trouble with Algebra 2 -_-, could someone maybe explain to me how I solve these problems?
(-3)^2 -2(6^2-9)
6-(4^2-7)^2
Thanks in advance =)

Answer by ptaylor(2198)

(-3)^2 -2(6^2-9)
Step 1: (-3)^2=9
Step 2:(6^2-9)=(36-9)=27
-2*27=-54
Step 3: Now put it back together: 9-54=-46

6-(4^2-7)^2
Step 1: (4^2-7)=(16-7)=9: 9^2=81
Step 2: Now put it back together: 6-81=-75
Get rid of the parens first
Thanks in advance =)

Hope this helps------------ptaylor